OpenStudy (anonymous):

\[\left| \theta \right|<\Pi/16 , then \sqrt{2 +\sqrt{2+\sqrt{2+ 2 \cos 8 teta}}}\] is equal to

OpenStudy (anonymous):

@satellite73 @apoorvk @ajprincess

OpenStudy (anonymous):

@shivam_bhalla

OpenStudy (anonymous):

the options are cos2theta 2costheta

OpenStudy (anonymous):

@supercrazy92 will help u surely :)

OpenStudy (anonymous):

lol, @heena is going to help you

OpenStudy (ajprincess):

2+2cos8x =2(1+cos8x) =2(1+2cos^2(4x)-1) =2*2c0s^2(4x) =4cos^2(4x) Find the root of it.

OpenStudy (anonymous):

no no its a golden opportunity for u to get a medal from me go for it :P @supercrazy92

OpenStudy (anonymous):

@ajprincess there are three 2 s ....

OpenStudy (ajprincess):

ya i knw that. did u find the root of it?

OpenStudy (anonymous):

2cos2theta

OpenStudy (ajprincess):

No it is 2cos4theta Nw do for 2+2cos4theta as I did for 2+2cos8theta.

OpenStudy (ajprincess):

2+2cos4theta =2(1+cos4theta) can u do nw?

OpenStudy (anonymous):

it seems finished

OpenStudy (anonymous):

i gave up!

OpenStudy (ajprincess):

show me yr work.

OpenStudy (anonymous):

i dont ghave any idea can u tell wat u have done

OpenStudy (ajprincess):

2+2cos8x =2(1+cos8x) (First I took 2 as the common factor) =2(1+2cos^2(4x)-1) (using cos2x=2cos^2x-1 I wrote this) =2(2cos^2(4x)) =4cos^2(4x) Is it clear nw?

OpenStudy (paxpolaris):

\[\large \cos \left( 2x \right)=\ \ \cos^2\left( x \right)-\sin^2\left( x \right)\]\[\large \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ =\ \cos^2\left( x \right)-\left( 1-\cos^2\left( x \right) \right)\]\[\LARGE \ \ \ \ \ \ \ \ \ \ =\ 2\cos^2\left( x \right)-1\] keep using this rule, as ajprincess showed, to show what's under a square root as a perfect square: then simplify from the inside out